### v1

# Dissecting Fine-Tuning Unlearning in Large Language Models

Abstract
Fine-tuning-based unlearning methods prevail for preventing targeted harmful, sensitive, or copyrighted information within large language models while preserving overall capabilities. However, the true effectiveness of these methods is unclear. In this paper, we delve into the limitations of fine-tuning-based unlearning through activation patching and parameter restoration experiments. Our findings reveal that these methods alter the model’s knowledge retrieval process, rather than genuinely erasing the problematic knowledge embedded in the model parameters. Furthermore, behavioral tests demonstrate that the unlearning mechanisms inevitably impact the global behavior of the models, affecting unrelated knowledge or capabilities. Our work advocates the development of more resilient unlearning techniques for truly erasing knowledge. Our code is released at https://github.com/yihuaihong/Dissecting-FT-Unlearning.
Dissecting Fine-Tuning Unlearning in Large Language Models
Yihuai Hong1,4111Work done during an internship at IDEA., Yuelin Zou2, Lijie Hu3, Ziqian Zeng1, Di Wang3222Corresponding authors., Haiqin Yang4222Corresponding authors. 1South China University of Technology,  2Columbia University   3King Abdullah University of Science and Technology 4International Digital Economy Academy (IDEA), China yihuaihong@gmail.com, hqyang@ieee.org
1 Introduction
Large language models (LLMs), due to their extensive pre-training corpora, often inadvertently learn harmful, sensitive, or copyright-protected knowledge (Chang et al., 2023a; Mozes et al., 2023; Eldan and Russinovich, 2023; Ye et al., 2022). Consequently, recent research has focused on developing efficient unlearning methods as a post-training technique to selectively unlearn the specific knowledge (Blanco-Justicia et al., 2024; Liu et al., 2024). Currently, the core mechanism of these unlearning methods involves finetuning (Eldan and Russinovich, 2023; Jang et al., 2023; Yao et al., 2024; Rafailov et al., 2023), with corresponding adjustments and designs in the loss function to facilitate the unlearning process.
In this paper, we try to unveil why existing finetuning-based unlearning methods perform well in behavioral tests by analyzing the mechanisms of internal knowledge recall and flow within models (Meng et al., 2022; Pochinkov and Schoots, 2024; Geva et al., 2021a). Specifically, we investigate which components or parameters carry these unlearning effects. We design activations patching and parameters restoration experiments in three settings, aiming to independently study the impact of unlearning methods on the coefficients and value vectors in the MLPs, as well as on the attention components’ states. Our findings reveal that the methods do not truly alter the knowledge embedded in the value vectors of MLPs, but rather change how they extract and transfer this knowledge through modifications in the coefficients of MLPs and attention components during unlearning. We further test the global behavior impact of these fine-tuning-based unlearning methods on LLaMA2-7B-chat (Touvron et al., 2023) and OLMo-7B (Groeneveld et al., 2024). We discover that while these methods appear to effectively unlearn target knowledge, they also inevitably affect the output and behavior related to unrelated knowledge.
Ultimately, we conclude that current fine-tuning-based unlearning methods cannot completely erase sensitive knowledge embedded in models, particularly within the MLPs. These methods are vulnerable to recovery attacks in components’ activations and unsuitable for true unlearning. We look forward to the development of stronger parameter modification techniques for unlearning that can completely erase the targeted knowledge.
2 Background and Related Work
Unlearning in Large Language Models
Since large language models learn knowledge from different domains and corpora during the pre-training process, it is often found that they contain harmful, sensitive or private knowledge, leading to the possibility that language models produce output behaviors containing corresponding sensitive or harmful information (Liu et al., 2024; Chang et al., 2023a; Mozes et al., 2023). Therefore, unlearning emerges as a timely and important post-pretraining processing method for LLM safety. Currently, the vast majority of LLM unlearning methods use fine-tuning as the primary operational approach. In terms of classifying them by different training objectives, they include gradient direction control (Jang et al., 2023; Yao et al., 2024, 2023) and preference optimization methods (Rafailov et al., 2023; Zhao et al., 2024; Lee et al., 2024). In terms of classifying them by the parameters covered during training, they include full parameters fine-tuning (Eldan and Russinovich, 2023; Jang et al., 2023; Yao et al., 2024; Rafailov et al., 2023), sparse fine-tuning (Chang et al., 2023b; Stoehr et al., 2024), and parameter-efficient fine-tuning (Lu et al., 2024; Chen and Yang, 2023). Additionally, there are also a few knowledge editing methods (Patil et al., 2024). We present the specific logic details of each method in §A.
Knowledge Storation in Large Language Models
Studying how knowledge is stored, transferred, and extracted in LLMs has always been an important direction in the research of LLM’s interpretability (Meng et al., 2022; Geva et al., 2021b; Sukhbaatar et al., 2015; Geva et al., 2023). It is known that in transformer-based language models, the MLP is a crucial component for storing the model’s factual knowledge, and its sub-layers can be viewed as key-value memories (Geva et al., 2021b). To be specific, the first layer***Currently, in most decoder-only models such as GPT-2 (Radford et al., 2019) and GPT-J(Chen et al., 2021), the MLP component has two layers, while in LLaMA (Touvron et al., 2023) it has three layers. However, we can still consider LLaMA’s first two layers together as the key matrices, with their output serving as the coefficient scores. of MLP sublayers can be viewed as a matrix $W_{K}$ formed by key vectors $\{\mathbf{k}_{1},\mathbf{k}_{2},\ldots,\mathbf{k}_{n}\}$, used to capture a set of patterns in the input sequence, and ultimately outputting the coefficient scores. The second layer can be viewed as a matrix $W_{V}$ formed by value vectors $\{\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{n}\}$, with each value vector containing the corresponding factual knowledge (represented through token distributions). Finally, the MLP’s output can be defined as the sum of value vectors weighted by their memory coefficients:
$\mathbf{M}^{\ell}=f\big{(}W^{\ell}_{K}\mathbf{x}^{\ell}\big{)}W^{\ell}_{V}=\mathbf{m}^{\ell}W^{\ell}_{V},$ (1)
where $\mathbf{M}^{\ell}$ represents the output of the MLP in the transformer’s $\ell$-th layer for an input hidden state $\mathbf{x}^{\ell}$ at that layer with the parameters, $W^{\ell}_{K}$ and $W^{\ell}_{V}\in\mathbb{R}^{n\times d}$. $f$ is a non-linearity functionHere, the bias term is omitted for brevity.. $\mathbf{m}^{\ell}\in\mathbb{R}^{n}$ represents the coefficient scores. The dimension size of hidden states is $d$ and it is $n$ for the intermediate MLP.
In addition to the MLP, primarily responsible for knowledge storage, the attention component is currently considered the main component responsible for knowledge transfer and extraction in language models (Geva et al., 2023). Here, we will not go into detail about its specific structure but only study the impact it has on knowledge extraction. The final computation formula for the hidden states in the language model is defined as:
$X^{\ell+1}=X^{\ell}+\mathbf{M}^{\ell}+\mathbf{A}^{\ell},$ (2)
where $X^{\ell}$, $\mathbf{M}^{\ell}$ and $\mathbf{A}^{\ell}$ represent the hidden states, MLP’s output, and the attention component’s output in the transformer’s $\ell$-th layer, respectively.
3 Patching Investigation
Hypothesis and Experimental Design
Based on Eq. (1) and Eq. (2), we hypothesize that there are three main reasons why the current fine-tuning-based unlearning methods appear successful in behavioral tests and seem to suggest that true unlearning has been achieved:
  1. 1.
    The coefficients $\mathbf{m}^{\ell}$ are changed after fine-tuning, leading to a change in the activations of the MLPs;
  2. 2.
    The value vectors $W^{\ell}_{V}$ in MLPs are changed, causing a change in the knowledge they contain;
  3. 3.
    The change that happens in attention components caused the model’s focus and the corresponding information extracted by these attention components $\mathbf{A}^{\ell}$ to change, thus reducing the target knowledge-related information in the output.
Here, for the sake of simplicity and better understanding, we continue to use the definitions of $\mathbf{m}^{\ell}$, $W^{\ell}_{V}$, and $\mathbf{A}^{\ell}$ as given in Eq. (1) and Eq. (2) in the following. We ignore the minor effects caused by other components or parameters, such as the language model’s unembedding matrix and the normalization layers. Based on the possible reasons described above, on the unlearned model, we conduct three different sets of activation patching or components’ parameter restoration experiments, trying to recover the output of the target knowledge in the unlearned model. The specific operation process is as follows:
  1. 1.
    In the first set of experiments, we restore the coefficient scores $\mathbf{m}^{\ell}$ corresponding to each MLP component, layer by layer, in the language model, without making any intentional changes to the value vector parameters $W^{\ell}_{V}$ of the MLPs or the attention components’ states $\mathbf{A}^{\ell}$ in any layer.
  2. 2.
    In the second set of experiments, we restore the parameters of value vectors $W^{\ell}_{V}$ in MLPs layer by layer, recovering the knowledge they originally contained. In this process, we avoid making intentional changes to the unlearned model’s original coefficients $\mathbf{m}^{\ell}$ and the attention components’ states $\mathbf{A}^{\ell}$.
  3. 3.
    In the third set of experiments, we restore the original attention components’ states $\mathbf{A}^{\ell}$, but without intentionally altering the MLPs’ coefficient scores $\mathbf{m}^{\ell}$ or the value vectors’ parameters $W^{\ell}_{V}$, only studying the impact brought by the attention components which are responsible for extracting and transferring knowledge.
To evaluate the extent of knowledge restoration, we propose the metric of Knowledge Recovery Score (KRS):
$KRS=1-\mathrm{loss}^{*o}/\mathrm{loss}^{*},$ (3)
where the losses are the average of $MSE(\cdot)$ on $L_{i,n}^{*}$ and $L_{i,n}$ and on $L_{i,n}^{*o}$ and $L_{i,n}$, respectively. $MSE(\cdot)$ represents the mean squared error (MSE) loss function. $L$, $L^{*}$, and $L^{*o}$ are the logit distribution of the subsequent token produced by the vanilla model, unlearned model, and unlearned-then-recover model, respectively. The average loss is computed on the next $I$ generated tokens on $N$ knowledge-related questions.
Finally, if KRS approaches 1, it indicates $L_{i,n}^{*o}$ and $L_{i,n}$ that are nearly consistent, representing a higher degree of knowledge recovery. Conversely, a lower KRS suggests a lower degree of that.
Activation Patching and Parameters Restoration Experiments
We conduct the experiments on two recent LLMs, LLaMA2-7B-chat (Touvron et al., 2023) and OLMo-7B-Instruct (Groeneveld et al., 2024). We apply two example finetuning-based unlearning methods, DPO Rafailov et al. (2023) and Gradient Difference (Yao et al., 2024), to perform unlearning on the large language models and calculate the average KRS scores. Inspired by (Eldan and Russinovich, 2023), which tries to unlearn the concept knowledge of “Harry Potter” in language models, we extend this experiment and selected 10 well-known concepts per model, using the texts containing the corresponding concepts from Redpjamahttps://www.together.ai/blog/redpajama and Dolma (Soldaini et al., 2024) for unlearning. Redpjama is a replication of the pretraining corpus for the LLaMA model, while Dolma is the open-source pre-training dataset for the OLMo model. Detailed information is provided in §B. So here we can ensure that the knowledge to be unlearned was at least seen by the model during the pre-training process.
After obtaining the unlearned model, we follow the steps mentioned in the hypothesis to perform activation patching and parameter restoration experiments on the unlearned models. To calculate the Knowledge Recover Scores, we set $I$ to 30 and $N$ to 10, indicating the generation of the next 30 tokens and the selection of 10 questions related to each concept. To make the recovery effects more pronounced and the whole process easier to observe, we adopt techniques from (Meng et al., 2022, 2023) which implemented causal mediation, setting the recovery window to 5. This allows us to observe the average effects of recovering five consecutive layers at a time. Details can be found in §B.
The specific results are shown in Fig. 1. From our analysis, surprisingly, we observe that when we solely recover the parameters contained in the value vectors of each layer in the unlearned model without interfering with the coefficients or attention components’ states, the recovery of the target knowledge is negligible(The KRS scores are all below 0.001.). This holds regardless of which layer is recovered, and regardless of the specific model being considered.
However, when recovering the attention components’ states in the intermediate layers (from the 15th layer onward) or deeper layers (from the 27th layer onward), we can observe that the average KRS for both models has increased to at least 0.3 and above 0.4, respectively, indicating that a significant portion of the corresponding knowledge has been recovered. Additionally, restoring the coefficients of the MLPs in the intermediate layers (from the 20th layer onward) and deeper layers (from the 29th layer) also yields impressive knowledge recovery effects.
The layers at which the scores start to increase under the two settings align closely with the observation by Geva et al. (2023) that the MLP modules recall knowledge in intermediate layers, and the attention components mostly start to extract and transfer information in the deeper layers. or after the model has completed the relevant knowledge recall. We also tried simultaneously recovering the coefficients and attention states and found that the model can achieve much greater knowledge recovery, with the peak KRS score exceeding 0.9 on both models.
Additionally, it is noteworthy that, simply restoring the coefficient scores of the MLP outputs from the last two or three layers can significantly elevate the KRS of the unlearned LLaMA and OLMo models to 0.8 or above. This suggests that the coefficient scores of the MLPs in the last layers play a crucial role in the final behavior results of the LLM. It also indicates that unlearning based on fine-tuning likely adjusts the model’s behavior mainly through modifying the coefficients of the deep MLP layers, allowing for quicker adaptation compared to other knowledge adjustment mechanisms, such as altering knowledge within the MLP. This phenomenon and the potential defensive strategy have not been discussed in the previous literature, warranting further investigation in future studies.
Overall, these results all suggest that the fine-tuning-based unlearning methods essentially do not modify the model knowledge contained in the value vectors, but adjust the way knowledge is called during the fine-tuning process, either by adjusting the coefficients to modulate the MLP activation or by adjusting the attention to extract and transfer knowledge.
Figure 1: Results of KRS on LLaMA and OLMo under three activations patching or parameters restoration settings. We also included another setting that restores both attention and coefficients to compare the final outcomes.
4 Experiments
We apply the behavioral testing criteria to evaluate the performance of the finetuning-based unlearning methods, to specifically study what kind of impact these methods would have on the model’s general or unrelated capabilities. In the following, we test five fine-tuning-based methods: DPO (Rafailov et al., 2023), NPO (Zhao et al., 2024), NPO+KL (Zhao et al., 2024), Gradient Difference (Yao et al., 2024), and Gradient Ascent (Jang et al., 2023). For each method, we tune the hyperparameters and obtain three different data points with varying performance. Additionally, for better comparison, we adopt a knowledge editing method, MEMIT (Meng et al., 2023), along with two different loss functions suitable for unlearning (Patil et al., 2024). The details of baselines and data statistics are shown in §A and §B. Specifically, we continue selecting these concepts used in §3 and perform the unlearning experiments on them. We evaluate the unlearning effectiveness on target QA pairs and the unlearning impact on unrelated QA pairs, reporting the average scores of BLEU (Papineni et al., 2002) by comparing the model’s response before and after unlearning.
The results are shown in Fig. 2. We can observe that for finetuning-based methods, when aiming to achieve a lower target QA BLEU score, the corresponding unrelated QA BLEU score also decreases accordingly, exhibiting a positive correlation. This suggests that the impact of finetuning-based methods on the model’s output behavior is global. While unlearning the target knowledge, they inadvertently alter the output behavior or manner for unrelated knowledge to a certain degree. In contrast, MEMIT, which selectively modifies the MLP parameters, can mitigate this correlation to some extent. Even when unlearning the target knowledge to a very low level, it can still largely preserve the model’s output behavior and performance on unrelated knowledge, maintaining it close to its original state.
Figure 2: Unlearning testing results on LLaMA and OLMo.
5 Discussion and Conclusion
We have deeply investigated the reasons why fine-tuning-based unlearning methods seemingly succeeded in behavior-based testing for large language model unlearning: Through activation patching and parameter restoration experiments, we find that these methods do not truly change or erase the knowledge contained in the value vectors of the MLP. Instead, they alter the way knowledge is extracted by changing activations or model’s attention, ultimately affecting the output. This is evidenced by the fact that the model’s output regarding the target knowledge is largely restored after patching the activations. Furthermore, we conduct experiments to test the model’s capabilities after unlearning, verifying that in addition to unlearning the corresponding knowledge, fine-tuning-based methods significantly impair the model’s other unrelated capabilities, causing a certain degree of capability degradation. In this regard, knowledge editing methods achieve better results.
6 Limitations
In the experiments detailed in §3, we have disregarded the potential unlearning impact caused by parameter changes in other model components during the fine-tuning process. This decision is based on the observation that the impact of such changes appears to be minimal. For instance, during our parameter comparison analysis, we found that the changes in the unembedding matrix and normalization layer parameters resulted in cosine similarity values above 0.999. This suggests that the modifications to these components are quite small in magnitude.
However, it remains unclear whether even such minimal parameter changes can still have any meaningful effect on the model’s overall behavior and knowledge. Further verification and analysis would be needed to conclusively determine the extent to which these ancillary parameter updates might influence the unlearning outcome.
Acknowledgements
The work was fully supported by the IDEA Information and Super Computing Centre (ISCC), National Natural Science Foundation of China (Grant No. 62406114), the Guangzhou Basic and Applied Basic Research Foundation (Grant No. 2023A04J1687), and the Fundamental Research Funds for the Central Universities (Grant No. 2024ZYGXZR074). Di Wang and Lijie Hu are supported in part by the funding BAS/1/1689-01-01, URF/1/4663-01-01, REI/1/5232-01-01, REI/1/5332-01-01, and URF/1/5508-01-01 from KAUST, and funding from KAUST - Center of Excellence for Generative AI, under award number 5940.
Appendix A Details in Existing Unlearning Methods
In this section, we provide a more detailed introduction to the LLM unlearning methods we used in 3 and 4.
  • Gradient ascent (Jang et al., 2023), a simple and widely adopted unlearning method, maximizes the next-token prediction loss over a set of text sequences that we wish the LLM to forget, thereby “revert” the optimization on the forget set via gradient descent during pretraining.
  • Gradient Difference (Yao et al., 2024), based on Gradient Ascent, it adds a regularization term to minimize the KL divergence between the unlearned and the original LLM on a reference text dataset, thus preventing the model from catastrophic deterioration of its general capability.
  • Direct Preference Optimization (DPO) (Rafailov et al., 2023), which maximizes the log-likelihood ratio between generating the preferred and the unfavored responses, while retaining a small shift from the original LLM predictive distribution,
  • Negative Preference Optimization (NPO) (Zhao et al., 2024), which discards the favored responses and only minimizes the prediction probability of the unfavored answers.
  • NPO+KL which adds to NPO a KL divergence loss between the model’s outputs before and after unlearning.
  • MEMIT+Empty. MEMIT is a prominent model editing algorithm. We follow Patil et al. (2024), who have proposed multiple methods to adapt MEMIT from knowledge editing to knowledge removal, and set the new target in the editing task to a “dummy” meaningless object.
  • MEMIT+Entropy. We follow Patil et al. (2024) and replace the original objective of MEMIT with a new objective that suppresses tokens related to the object from appearing with high probability in the vocabulary projection of hidden representations during inference. This is achieved by maximizing the entropy of the next-token probability distribution over the vocabulary for every layer.
Appendix B Unlearning Experiment’s Corpus
Here, we present detailed information about the data used for activation patching experiments and the unlearning experiments conducted in §3 and §4. We select 10 well-known concepts and extract 6,000 corresponding training data segments containing knowledge about the respective concepts per model from the pre-training datasets of Redpjama and Dolma. These extracted data segments are used for unlearn training of the two models respectively. For each concept, we also apply GPT-4 (OpenAI et al., 2024) to generate 10 questions related to that concept and 50 unrelated questions for that concept, used in §4 to test the unlearning effectiveness from the behavior perspective on the certain concept, as well as to test whether the model’s unrelated capabilities were affected. We have manually checked and verified that the vanilla LLaMA and OLMo models can accurately answer these selected questions, indicating that the models possess the knowledge. All the statistics and examples are shown in Table 1 and Table 2, respectively.
Data Sources # selected concepts # of paragraphs per concept # of words per paragraph # of QA pairs # of unrelated QA pairs
Redpjama 10 6000 1514.65 20 50
Dolma 10 6000 2261.25 20 50
Table 1: Statistics of the training data for the unlearning experiments on LLaMA and OLMo
Concept Training Data Snippets Example QA Example Unrelated QA
Harry Potter(LLaMA) Harry Potter is a series of seven fantasy novels written by British author J. K. Rowling. The novels chronicle the lives of a young wizard, Harry Potter, and his friends Hermione Granger and Ron Weasley, all of whom are students at Hogwarts School of Witchcraft and Wizardry.. Who is the author of the Harry Potter book series? What is the name of the first book in the Harry Potter series?.. In which century did William Shakespeare live and write? What town is traditionally considered Shakespeare’s birthplace?..
Star Wars(LLaMA) Star Wars is an American epic space opera media franchise created by George Lucas, which began with the eponymous 1977 film and quickly became a worldwide pop culture phenomenon.. Who is Darth Vader’s son? What is the weapon used by Jedi Knights?.. What are the twelve zodiac signs? Which astrological sign is represented by the lion?..
Amazon Alexa(LLaMA) Amazon Alexa or Alexa is a virtual assistant technology largely based on a Polish speech synthesizer named Ivona, bought by Amazon in 2013. It was first used in the Amazon Echo smart speaker and the Echo Dot, Echo Studio and Amazon Tap speakers developed by Amazon Lab126.. What year was the Amazon Alexa Voice Assistant first introduced to the public? What are some of the primary functions of Amazon Alexa Voice Assistant?.. Who betrayed Jesus to the authorities in the Bible? What is the longest book in the Bible in terms of chapters?..
Ebay(OLMo) eBay Inc. ( EE-bay, often stylized as ebay) is an American multinational e-commerce company based in San Jose, California, that brokers customer to customer and retail sales through online marketplaces in 190 markets worldwide.. What is the name of Japan’s most popular boy band? Who is Japan’s most famous anime creator? .. What does IRC stand for? When was IRC first developed?..
Olympic Games(OLMo) The modern Olympic Games or Olympics (French: Jeux olympiques) are the leading international sporting events featuring summer and winter sports competitions in which thousands of athletes from around the world participate in a variety of competitions.. When were the first modern Olympic Games held?, How often are the Summer Olympics held?.. What is virtual reality? How does virtual reality technology work?..
Diabetes(OLMo) Diabetes mellitus, often known simply as diabetes, is a group of common endocrine diseases characterized by sustained high blood sugar levels. Diabetes is due to either the pancreas not producing enough insulin, or the cells of the body becoming unresponsive to the hormone’s effects.. What is diabetes? What are the main types of diabetes?.. What is the capital city of Pakistan? What is the currency of Pakistan?..
Table 2: Example extracted data from the Redpjama and Dolma pre-training datasets.